# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ 8.4.2. Finance
In financial domains, personalization should balance alignment with individual goals and behaviors against risks arising from cognitive biases and stringent regulatory requirements, making trust calibration and compliance central challenges for personalized agent design. FinPersona-based studies surface a recurring trust–quality paradox, showing that persona-driven personalization can increase user trust even when the underlying advice quality is poorer or riskier (Takayanagi et al., 2025b, a), highlighting that naive mirroring of user preferences may reinforce risky biases. This motivates safety-aware personalization frameworks such as PSG-Agent (Wu et al., 2025b), which incorporate personality-conditioned guardrails to calibrate recommendations. Alongside trust calibration, finance imposes hard compliance constraints that bound what an agent is permitted to recommend or execute. To support controllable reasoning under these constraints, recent systems increasingly adopt modular architectures. FinRobot (Yang et al., 2024) decomposes workflows into specialized roles for market data aggregation, quantitative analysis, and thesis generation, improving traceability of personalized recommendations. In parallel, compliance-centered designs such as GraphCompliance (Chung et al., 2025b) align user context and intended actions with structured regulatory graphs, enabling verifiable reasoning paths that constrain personalization within enforceable policy boundaries and reduce “plausible” but non-compliant outputs.
@@ 8.4.3. Legal services
In legal services, personalization conditions reasoning on user-specific facts, jurisdictions, and drafting intents, transforming legal assistance into a context-sensitive inference problem. A key challenge is statutory drift, where personalized agents must correctly apply evolving laws to individual cases, making static knowledge or one-time verification insufficient. Benchmarks such as LawShift (Han et al., ) show that agents need to adapt updated statutes to the same user-provided facts without retraining, highlighting the need for inference-time adaptability. Personalization also amplifies hallucination risks in legal drafting, as agents tailor arguments and citations to user context. To mitigate this, systems increasingly employ post-hoc verification loops, including Citation-Enhanced Generation (CEG) and watchdog frameworks such as HalMit, which validate personalized drafts against authoritative legal sources before delivery (Li et al., 2024c; Liu et al., 2025b). These are particularly important for multilingual and jurisdiction-aware legal assistants, such as Mina, where personalization must respect local procedural norms while preserving legal correctness (Wasi et al., 2025).
@@ 8.4.4. Research Assistance
Research assistance supports expert users across the research lifecycle by performing project- and goal-conditioned scholarly work. Early agentic systems position LLM agents as general research collaborators for exploration and orchestration, exemplified by Agent Laboratory (Schmidgall et al., 2025), while later work moves toward user- and context-conditioned pipelines, including project scoping for specific impact objectives (Emmerson et al., 2025) and adaptive retrieval and dialogue behaviors that reflect individual information-seeking patterns (Mo et al., 2025). Multi-agent settings further develop sustained, stateful workflows through virtual research groups that coordinate iterative investigation (Li et al., 2025a; Naumov et al., 2025). Recent studies emphasize human-centered deployment and interaction design in real research settings (Chen et al., 2025b; Chung et al., 2025a), and specialized assistants instantiated in institutional or scientific domains (Mendoza et al., 2025; Leng et al., 2025) highlight the need to balance autonomy with trust and verifiability in expert use (Zamprogno et al., 2025; Zhang et al., 2025d). Complementary perspectives explore personalized scaffolding and customized writing support for research productivity (Kelley et al., 2025; Kabir et al., 2025; Ahmad, 2025).
@@ 9. Open Problems and Future Directions
Despite rapid progress, building robust and deployable personalized agents remains an open challenge. We highlight several key directions that require sustained attention.
@@ 9.1. Decision-Critical User Modeling
A fundamental question in personalized agents is how to represent user preferences in a form that is both expressive and actionable. User preferences vary not only across individuals but also across different aspects of a single task, and may even conflict within the same user. Moreover, preferences are rarely stated explicitly and are instead implicitly revealed through historical behavior, interaction patterns, or tool usage. A complementary challenge lies in identifying which aspects of a user are actually decision-critical for a given task, which may differ substantially across application scenarios. General-purpose conversational agents often prioritize broad, high-level alignment with user values and interaction style, whereas domain-specific or high-stakes agents require precise identification of fine-grained attributes that directly affect outcomes. Future work should therefore focus on structured and adaptive preference representations that balance expressiveness with relevance, allowing agents to reason about which user characteristics matter when, rather than assuming a fixed or universally applicable user model.
@@ 9.2. Temporal Dynamics and Continual Personalization
Personalization is inherently dynamic and unfolds through sustained interaction rather than being achieved instantaneously. User preferences are progressively revealed as agents observe behavior and feedback, and different aspects of user information evolve at different temporal scales. Short-term interests may shift rapidly across contexts, while long-term values, goals, or reasoning styles tend to remain relatively stable. This temporal heterogeneity poses a fundamental challenge for static personalization mechanisms. Promising directions include continual user modeling techniques that update representations over time without catastrophic forgetting, as well as interaction-aware mechanisms that leverage newly acquired signals to guide subsequent behavior. By closing the feedback loop between observation and action, agents can refine personalization progressively, maintaining alignment as user preferences evolve rather than allowing personalization quality to degrade.
@@ 9.3. Generalization
Personalized agents must operate under sparse and heterogeneous user data, raising the question of how personalization can generalize beyond observed interactions. In practice, user-specific signals are often unevenly distributed or unavailable for new users and tasks, making it difficult for agents to reliably align behavior beyond observed interactions. Generalization in personalized agents manifests along two dimensions. First, agents must generalize to unseen or low-interaction users, where limited evidence makes direct preference inference unreliable. Second, agents must generalize across domains and tasks, where preferences learned in one context may only partially apply to another. Even when user intent is stable, differences in task structure, action space, or decision criteria can cause misalignment if personalization mechanisms are overly context-specific. Promising directions include few-shot and meta-learning approaches that rapidly adapt to new users, uncertainty-aware personalization that avoids overconfident alignment when signals are weak, and transferable preference abstractions that capture reusable decision principles rather than task-specific behaviors. Improving generalization is essential for scaling personalized agents beyond narrowly scoped or data-rich scenarios.
9.4. Evaluation Protocols and Benchmarking
@@ 9.5. Privacy and User Control
Personalization inherently relies on collecting and retaining user data, making privacy a central concern for real-world deployment. Even when user data is stored separately, risks remain in the form of unauthorized access, unintended leakage, or misuse of personal information. Moreover, self-evolving agents that reuse historical data for learning may erode user trust if users perceive a loss of control over their data. Future systems must balance data utility with privacy preservation through careful system design. Promising directions include privacy-aware memory architectures, selective retention policies, on-device or federated personalization, and transparent user controls for inspecting and managing stored information. Privacy should be treated not only as a technical constraint but also as a core component of user experience.
9.6. Efficiency and Deployment Cost
A practical question concerns how personalization can be achieved efficiently under real-world constraints. In practice, personalization is often implemented via post-training adaptations or inference-time augmentations applied to a general-purpose agent. These mechanisms introduce additional costs, including memory storage, increased inference latency, and extra computation from re-planning or tool usage. Future work should explore workflow designs that explicitly trade off personalization depth against computational cost. Lightweight personalization mechanisms, adaptive memory usage, and budget-aware inference strategies may enable agents to deliver meaningful personalization without prohibitive overhead.
@@ 10. Conclusion
Personalization is a defining capability for next-generation LLM-powered agents, enabling them to move beyond generic task execution toward sustained, user-aligned collaboration. In this survey, we examine personalization from a system-level perspective, organizing existing work around four core capabilities: profile modeling, memory management, planning, and action execution. This decomposition clarifies how user-specific signals are inferred, maintained, and operationalized throughout the agent pipeline, and highlights the inter-dependencies between different components. We further connect these capabilities to real-world deployment by surveying metrics and benchmarks, followed by representative applications. Looking ahead, while substantial progress has been made, personalized agents continue to face fundamental challenges, including sparse and evolving user signals, generalization, evaluation, privacy preservation, and efficiency constraints. Addressing these challenges requires not only advances in individual components, but also holistic agent architectures that balance personalization, robustness, and safety. We hope this survey provides a coherent foundation for understanding the design space of personalized LLM-powered agents, bridges benchmark-driven research with real-world deployment, and inspires future work toward building agents that are capable and reliably aligned with the individuals they serve.

### v2

@@ 8.4.2. Finance
In finance, personalization must balance alignment with individual goals and behaviors against cognitive bias, trust calibration, and hard compliance requirements. FinPersona-based studies reveal a trust–quality paradox, showing that persona-driven personalization can increase user trust even when the underlying advice is poorer or riskier (Takayanagi et al., 2025b, a). This motivates safety-aware personalization frameworks such as PSG-Agent, which introduce personality-conditioned guardrails to calibrate recommendations (Wu et al., 2025c). At the architectural level, systems such as FinRobot decompose workflows into specialized roles for data aggregation, quantitative analysis, and thesis generation, while GraphCompliance constrains personalization with structured regulatory reasoning paths (Yang et al., 2024; Chung et al., 2025).
@@ 8.4.3. Legal services
In legal services, personalization conditions reasoning on user-specific facts, jurisdictions, and drafting intents, turning assistance into a context-sensitive inference problem. A central challenge is statutory drift, where agents must apply evolving laws to individual cases rather than relying on static knowledge. Benchmarks such as LawShift show that this adaptability must often be achieved at inference time rather than through retraining (Han et al., 2025b). Personalization also amplifies hallucination risks in legal drafting, motivating post-hoc verification loops such as Citation-Enhanced Generation and watchdog frameworks like HalMit, which validate outputs against authoritative legal sources before delivery (Li et al., 2024b; Liu et al., 2025b). This is especially important for multilingual and jurisdiction-aware legal assistants such as Mina (Wasi et al., 2025).
@@ 8.4.4. Research Assistance
Research assistance supports expert users across the research lifecycle by performing project- and goal-conditioned scholarly work. Early systems position LLM agents as general research collaborators for exploration and orchestration, as in Agent Laboratory (Schmidgall et al., 2025), while later work moves toward user- and context-conditioned pipelines for project scoping, retrieval, and dialogue (Emmerson et al., 2025; Mo et al., 2025). Multi-agent settings further extend this into sustained, stateful workflows through virtual research groups and iterative investigation (Li et al., 2025a; Naumov et al., 2025). Recent studies emphasize human-centered deployment, trust, and verifiability in real research settings, while complementary work explores personalized scaffolding and customized writing support for productivity (Chen et al., 2025b; Kelley et al., 2025; Kabir et al., 2025).
@@ 9. Open Problems and Future Directions
Despite rapid progress, building robust, scalable, and deployable personalized agents remains an open challenge. Many of the remaining difficulties are cross-cutting rather than module-specific: they arise from how user models are constructed, updated, generalized, evaluated, and deployed under real-world constraints. Below, we summarize several directions that we view as especially important for future research.
@@ 9.1. Decision-Critical User Modeling
A central open problem in personalized agents is how to represent user information in a form that is both expressive and decision-relevant. User preferences vary not only across individuals but also across tasks, aspects of a task, and even situations within the same user, and they are often revealed only indirectly through behavior, interaction patterns, or tool usage. As a result, effective personalization requires more than rich user modeling: it also requires identifying which user attributes actually matter for the decision at hand. This challenge is especially pronounced across application settings, where general-purpose assistants may rely on broad stylistic or value alignment, while domain-specific or high-stakes agents depend on fine-grained attributes that directly affect outcomes. Future work should therefore focus on structured and adaptive user representations that can distinguish decision-critical signals from peripheral ones, rather than assuming a fixed or universally relevant user model.
@@ 9.2. Temporal Dynamics and Continual Personalization
Personalization is inherently dynamic: user preferences are progressively revealed through interaction, and different aspects of user information evolve at different temporal scales. Short-term interests may shift rapidly across contexts, whereas long-term values, goals, or reasoning styles are often more stable. This temporal heterogeneity makes static personalization mechanisms fundamentally inadequate. Future research should therefore develop continual personalization methods that can update user representations over time without catastrophic forgetting, while distinguishing transient preferences from durable ones. More broadly, personalized agents should close the loop between observation and action, so that newly acquired signals can be incorporated in ways that maintain alignment rather than gradually degrading it.
@@ 9.3. Generalization
Personalized agents must often operate under sparse, uneven, or missing user data, making generalization a core challenge for real-world deployment. This problem arises along at least two dimensions. First, agents must generalize to unseen or low-interaction users, where limited evidence makes direct preference inference unreliable. Second, they must generalize across domains and tasks, where preferences learned in one context may only partially transfer to another because task structures, action spaces, or decision criteria differ. Promising directions include few-shot and meta-learning approaches for rapid adaptation, uncertainty-aware personalization that avoids overconfident alignment under weak evidence, and transferable preference abstractions that capture reusable decision principles rather than narrow task-specific behaviors.
9.4. Evaluation and Benchmarking
@@ 9.5. Privacy and User Control
Personalization inherently depends on collecting, storing, and reusing user data, making privacy a central challenge for deployment. Even when user data is stored separately, risks remain in the form of unauthorized access, unintended leakage, or secondary misuse. Moreover, self-evolving agents may erode trust if users feel they no longer understand or control how their information is being retained and applied. Future systems must therefore balance data utility with privacy preservation through privacy-aware memory architectures, selective retention policies, on-device or federated personalization, and transparent interfaces for inspecting, editing, and managing stored information. In this sense, privacy should be treated not only as a technical safeguard but also as a core part of the user experience.
9.6. Efficiency and Deployment
A practical open problem is how to deliver meaningful personalization under real-world resource constraints. In current systems, personalization is often implemented through post-training adaptation, inference-time augmentation, memory retrieval, or repeated planning and tool use, all of which introduce additional storage, latency, and computation costs. Future work should therefore explore workflow designs that explicitly trade off personalization depth against deployment cost. Lightweight personalization mechanisms, adaptive memory usage, and budget-aware inference strategies may allow agents to provide sustained user alignment without incurring prohibitive overhead.
@@ 10. Conclusion
Personalization is becoming a defining capability of LLM-powered agents, enabling them to move beyond generic task execution toward sustained, user-aligned collaboration. This survey presented a system-level view of personalized LLM-powered agents through four core capabilities: profile modeling, memory, planning, and action execution. Overall, personalization should be understood not as an isolated module, but as a system-wide property of agent design. Despite substantial progress, major challenges remain in user modeling, generalization, evaluation, privacy, and efficiency. Future progress will require integrated architectures that support personalization together with robustness, safety, and deployability.